If u multiply a number by five and then subtract negative ten,the difference is negative 35. What is the number?
Answer by macston(5194)

5x-(-10)=-35 add (-10) to each side
5x-(-10)+(-10)=-35+(-10)
5x=-45
x=-9
Negative 9 times 5 -(-10) equals -35